# NP-Hardness and ETH-Based Inapproximability of Communication Complexity via Relaxed Interlacing

### v1

@@ 4.2 Construction
Definition 12 (Transpose).
The transpose of an $m\times n$ Boolean matrix $M$ is the matrix $M^{\top}\in\{0,1\}^{n\times m}$ defined by $M^{\top}(i,j)=M(j,i)$ for all $i\in[n]$ and $j\in[m]$.
@@ 4.3 Correctness Proof
Table 3 gives a table of the parameters used to instantiate the reduction.
@@
Any matrix $N$ in $\mathcal{M}_{1}^{\top}\left[1,\left\lceil q_{2}\right\rceil_{\!2},\left((1+\varepsilon_{q_{2},t_{2}})\,h^{\prime}(t_{2})\,\left\lceil q_{2}\right\rceil_{\!2}-\varepsilon_{q_{2},t_{2}}\right)^{1/t_{2}}\right]$ is $(\delta,2)$-robust.
@@
Any matrix $N$ in $\mathcal{M}_{2}^{\top}\left[1,\left\lceil q_{3}\right\rceil_{\!2},\left((1+\varepsilon_{q_{3},t_{3}})\,h^{\prime}(t_{3})\,\left\lceil q_{3}\right\rceil_{\!2}-\varepsilon_{q_{3},t_{3}}\right)^{1/t_{3}}\right]$ is $(\delta,2)$-robust.
@@
Lemma 19 gives us the main condition (9) needed to apply Theorem 2. The proof follows the same arguments as 20 with the parameters one level down. ∎
@@
$\mathcal{M}_{2}\;=\;\mathcal{M}_{1}^{\top}\left\langle q_{2},\,S_{q_{2},t_{2}}(C_{1})\right\rangle,\qquad R_{2}=[q_{2}]\times C_{1},\quad C_{2}:=S_{q_{2},t_{2}}(C_{1}),$
@@
$\mathcal{M}_{1}^{(i)}\;:=\;\mathcal{M}_{2}\!\left[\,\{i\}\times C_{1},\,C_{2}\right],\qquad R_{i}:=\{i\}\times C_{1}.$
(Multiplicity lower bound). By the almost–kk–wise independence of the reservoir $S_{q_{2},t_{2}}(C_{1})$ we have, for every $c\in C_{1}$,
$\left\{\,\mathbf{d}\in C_{2}:d_{i}=c\right\}\;\;\geq\;\;\left(\tfrac{1}{|C_{1}|}-\varepsilon\right)\,|C_{2}|\;=:\;\mu.$
@@
$|C^{\prime}|\;\;\geq\;\;|C_{2}|-\mu\;=\;\left(1-\tfrac{1}{|C_{1}|}+\varepsilon\right)\,|C_{2}|.$
@@
In particular, the set of genuine rows of every embedded $\mathcal{M}_{1}^{(i)}$ is fully contained in $\mathcal{M}_{2}[R_{i},C^{\prime}]$; no authentic row is lost. We therefore say that each $\mathcal{M}_{1}$ inside $\mathcal{M}_{2}$ is complete.
@@
$C_{2}^{(i,c)}\;:=\;\{\mathbf{d}\in C_{2}\mid d_{i}=c\},\qquad c\in C_{1}.$
1. A multiplicity lower bound.
Because $C_{2}=S_{q_{2},t_{2}}(C_{1})$ is built with accuracy parameter $\varepsilon=\varepsilon_{q_{2},t_{2}}$, every pattern in any coordinate is $\varepsilon$–close to uniform. Hence for every $c\in C_{1}$ (and for every $i\in[q_{2}]$)
$|C_{2}^{(i,c)}|\;\;\geq\;\;\left(\tfrac{1}{|C_{1}|}-\varepsilon\right)\,|C_{2}|\;=:\;\mu.$
2. Choosing the fullness threshold.
Let $C^{\prime}\subseteq C_{2}$ be such that
$|C^{\prime}|\;\;\geq\;\;|C_{2}|-\mu\;=\;\left(1-\tfrac{1}{|C_{1}|}+\varepsilon\right)\,|C_{2}|.$
Equivalently, fewer than $\mu$ columns of $C_{2}$ are deleted.
3. No column of $\boldsymbol{C_{1}}$ disappears.
Take any $c\in C_{1}$. From ($\star$) there are at least $\mu$ columns in $C_{2}$ whose $i$-th coordinate equals $c$. Since at most $\mu-1$ columns are dropped when forming $C^{\prime}$, at least one column survives with $d_{i}=c$. Thus $C^{\prime}\cap C_{2}^{(i,c)}\neq\varnothing$ for every $c\in C_{1}$.
[…]
Row $(i,c)$ of $\mathcal{M}_{2}$ depends only on the $i$-th coordinate of the column. Because for every $c\in C_{1}$ some column with $d_{i}=c$ remains in $C^{\prime}$, the entire set of distinct rows $\{(i,c)\mid c\in C_{1}\}$ is contained in $\mathcal{M}_{2}[R_{i},C^{\prime}]$. What may be missing are duplicates: if several columns share the same pair $(i,c)$, discarding a subset of them deletes only redundant copies of that row. Duplicates are irrelevant for deterministic communication complexity, so we may identify $\mathcal{M}_{2}[R_{i},C^{\prime}]$ with $\mathcal{M}_{1}^{(i)}$ “up to the deletion of duplicated rows”.
5. Completeness.
Since the argument is uniform in $i$, every embedded $\mathcal{M}_{1}^{(i)}$ is preserved inside $\mathcal{M}_{2}[R_{i},C^{\prime}]$ in the above sense. Hence each copy of $\mathcal{M}_{1}$ inside $\mathcal{M}_{2}$ is complete, finishing the proof. ∎
@@ 6.1 Relationship to Concurrent Work
While this manuscript was being written, Hirahara et al. [3] independently showed NP-hardness of deciding $D(f)\geq k$ using an unbounded number of alternations, which they highlight as an inherent limitation of their approach. Our constant-alternation hardness resolves this issue: it rules out the possibility (assuming $P\neq NP$) that any function $f$ with bounded-round optimal protocol admits an efficient exact algorithms to compute $CC(f)$ even if more general functions do not. Note that this is distinct from the result by Hirahara et al. [2] which shows that if the protocol is only allowed to alternate a constant number of times, deciding communication complexity is NP-hard. Our reduction shows NP-hardness for functions that can be solved optimally in constant rounds.
Moreover, our ETH-based additive inapproximability work sketches out how to show that approximating $D(f)$ within $\alpha(n)\in\omega(1)$ is hard under standard fine-grained assumptions, a direction also mentioned as wide-open in the concurrent work.

### v2

@@ 4.3 Correctness Proof
Table 3 gives a table of the parameters used to instantiate the reduction.11todo: 1add $d$ and comment it is power of 2, Notation for power of 2
@@
Any matrix $N$ in $\mathcal{M}_{1}^{\top}\left[1,\left\lceil q_{2}\right\rceil_{\!2},\left(h^{\prime}(t_{2})\,\left\lceil q_{2}\right\rceil_{\!2}/(1+\varepsilon_{q_{2},t_{2}})\right)^{1/t_{2}}\right]$ is $(\delta,2)$-robust.
@@
Any matrix $N$ in $\mathcal{M}_{2}^{\top}\left[1,\left\lceil q_{3}\right\rceil_{\!2},\left(h^{\prime}(t_{3})\,\left\lceil q_{3}\right\rceil_{\!2}/(1+\varepsilon_{q_{3},t_{3}})\right)^{1/t_{3}}\right]$ is $(\delta,2)$-robust.
@@
Lemma 18 gives us the main condition (11) needed to apply Theorem 2. The proof follows the same arguments as 19 with the parameters one level down. ∎
@@
$\mathcal{M}_{2}\;=\;\mathcal{M}_{1}^{\top}\left\langle q_{2},\,S_{q_{2},t_{2}}(R_{1})\right\rangle,\qquad R_{2}=[q_{2}]\times C_{1},\quad C_{2}:=S_{q_{2},t_{2}}(R_{1}),$
@@
$\mathcal{M}_{1}^{(i)}\;:=\;\varepsilon\left(\mathcal{M}_{2};\{i\}\times C_{1},C_{2}\right)\qquad R_{i}:=\{i\}\times C_{1}.$
(Multiplicity lower bound). By the almost–kk–wise independence of the reservoir $S_{q_{2},t_{2}}(C_{1})$ we have, for every $c\in R_{1}$,
$|\left\{\mathbf{d}\in C_{2}\mid d_{i}=c\right\}|\;\;\geq\;\;\tfrac{1-\varepsilon}{|R_{1}|}\cdot|C_{2}|\;=:\;\mu.$
@@
$|C^{\prime}|\;\;\geq\;\;|C_{2}|-\mu+1\;=\;\left(1-\tfrac{1-\varepsilon}{|R_{1}|}\right)\,|C_{2}|.$
@@
In particular, the set of genuine rows of every embedded $\mathcal{M}_{1}^{(i)}$ is fully contained in $\varepsilon\left(\mathcal{M}_{2};R_{i},C^{\prime}\right)$; no authentic row is lost. We therefore say that each $\mathcal{M}_{1}$ inside $\mathcal{M}_{2}$ is complete.
@@
$C_{2}^{(i,c)}\;:=\;\{\mathbf{d}\in C_{2}\mid d_{i}=c\},\qquad c\in R_{1}.$
No column of $\boldsymbol{R_{1}}$ disappears.
Take any $c\in R_{1}$. There are at least $\mu\leq C_{2}^{(i,c)}$ columns in $C_{2}$ whose $i$-th coordinate equals $c$. Since at most $\mu-1$ columns are dropped when forming $C^{\prime}$, at least one column survives with $d_{i}=c$. Thus $C^{\prime}\cap C_{2}^{(i,c)}\neq\varnothing$ for every $c\in R_{1}$.
Since the argument is uniform in $i$, every embedded $\mathcal{M}_{1}^{(i)}$ is preserved inside $\varepsilon\left(\mathcal{M}_{2};R_{i},C^{\prime}\right)$ in the above sense. Hence each copy of $\mathcal{M}_{1}$ inside $\mathcal{M}_{2}$ is complete, finishing the proof. ∎
@@ 6.1 Relationship to Concurrent Work
While this manuscript was being written, Hirahara et al. [2] independently showed NP-hardness of deciding $D(f)\geq k$. The two approaches use different techniques. Whilst their work gives a simpler reduction, the structural lemmas derived in this paper give us greater flexibility in the design of reductions. This opens up the exploration of reductions from problems where an inapproximability gap exists. This is highlighted in Hirahara et al. [2] as a limitation of their approach, and we believe our framework opens up possibilities in that direction.
Moreover, our ETH-based additive inapproximability work sketches out how to show that approximating $D(f)$ within $\alpha(n)\in\omega(1)$ is hard under standard fine-grained assumptions. This is possible thanks to the self-similarity inherent in our construction, and with further refinements it should be possible to extend inapproximability to stronger conditions when restricted to constants (i.e. NP-hardness of constant gap approximation).